# Solving rescheduling problems in heterogeneous urban railway networks using hybrid quantum-classical approach

### v4

@@ Acknowledgments
This research was supported by the Ministry of Culture and Innovation and the National Research, Development and Innovation Office within the Quantum Information National Laboratory of Hungary (Grant No. 2022-2.1.1-NL-2022-00004) (MK), and by the Silesian University of Technology Rector’s Grant no. BKM - 720/RT2/2023 12/020/BKM_2023/0252 (KK). This research was funded in part by National Science Center, Poland, under grant number 2019/33/B/ST6/0201 (LB) and 2023/07/X/ST6/00396 (KD). For the purpose of Open Access, the authors has applied a CC-BY public copyright license to any Author Accepted Manuscript (AAM) version arising from this submission.
@@ Appendix A. Train diagrams of CPLEX and CQM
To further evaluate the solutions, we present time-distance diagrams for network $7$, for the CPLEX solution and two different quantum realizations; these can be seen in Fig. 6 (left CPLEX, middle and right two realizations of CQM). We have chosen the line segment between GLC and KZ, offering a good demonstration. (We have introduced a small artificial ’distance’ between KO and KO(STM) which are contiguous locations for sake of better visibility, hence the horizontal lines in the diagram.) The red dotted lines represent the conflicted situation, and the green solid lines represent the solution. All the three proposed solutions have similar figures of merit, they all provide a valid option to be realized. The conflicts are resolved in somewhat different ways, see e.g. the paths of trains 40150, 5312, and 26103. In details:
@@
This results in the following objective ×$\times d_{\max}$ values: for CPLEX - $188.75$, for CQM first realization - $279$ for CQM second realization - $261.5$. Although the first solution is optimal, it is reasonable to supply dispatchers with the range of feasible solutions. Hence, these diagrams proved that the CQM hybrid solver can produce usable solutions for railway traffic management.
An additional benefit of the probabilistic nature of the quantum-based method is that it produces different alternatives after each realization run without any modification. These could be offered to the dispatcher in a decision support system as possibilities. The dispatcher’s choice can be influenced by other circumstances not covered by the present model.
Figure 6: Real-life experiment. Time-distance diagram solution of network $7$ CPLEX (left) and two realizations of CQM (middle and right), see Tab. 2. Real-life experiment. Time-distance diagram CQM solution of network $7$ (see Tab. 2) first realization. Closures (yielding single-track line) are between CB and KTC as well as between KO(STM) and KZ.

### v5

@@ Acknowledgments
This research was supported by the Ministry of Culture and Innovation and the National Research, Development and Innovation Office within the Quantum Information National Laboratory of Hungary (Grant No. 2022-2.1.1-NL-2022-00004) (MK), and by the Silesian University of Technology Rector’s Grant no. BKM - 720/RT2/2023 12/020/BKM_2023/0252 (KK). This research was funded in part by National Science Center, Poland, under grant number 2019/33/B/ST6/02011 (LB) and 2023/07/X/ST6/00396 (KD). For the purpose of Open Access, the authors have applied a CC-BY public copyright license to any Author Accepted Manuscript (AAM) version arising from this submission.
@@ Appendix A. Train diagrams of CPLEX and CQM
To further evaluate the solutions, we present time-distance diagrams for network $7$, for the CPLEX optimal solution and two different quantum realizations; these can be seen in Figure 6 (left: exact CPLEX, middle and right: two realizations from CQM). We have chosen the line segment between GLC and KZ, offering a good demonstration. (We have introduced a small artificial ’distance’ between KO and KO(STM) which are contiguous locations for the sake of better visibility, hence the horizontal lines in the diagram.) The red dotted lines represent the conflicted situation, and the green solid lines represent the solution. All three proposed solutions have similar figures of merit, they all provide a valid option to be realized. The conflicts are resolved in somewhat different ways, see e.g. the paths of trains 40150, 5312, and 26103. In details:
@@
This results in the following objective values: for CPLEX - $188.75$, for CQM first realization - $279$ for CQM second realization - $261.5$. Although the first solution is optimal, it is reasonable to supply dispatchers with the range of feasible solutions. These diagrams illustrate that the CQM hybrid solver can produce usable solutions for railway traffic management. An additional benefit of the probabilistic nature of the quantum-based method is that it produces different alternatives after each sampling, without any modification. These could be offered to the dispatcher in a decision support system as possibilities. The dispatcher’s choice can be influenced by other circumstances not covered by the present model.
Figure 6: Real-life experiment. Time-distance diagram solution of network $7$ CPLEX (left) and two realizations of CQM (middle and right), see Table 6. Real-life experiment. Time-distance diagram CQM solution of network $7$ (see Table 6) first realization. Closures (yielding single-track line) are between CB and KTC as well as between KO(STM) and KZ.